If the central angle =10 rad, and the area of the sector formed is 25 square units, find the area of the circle.
1
Expert's answer
2013-07-12T08:55:25-0400
This sector area is S=pi*r^2*(10-2pi)/2pi=pi*r^2*0.592=25 So area of the circle is pi*r^2=42.26 square units.